# Estimating Heterogeneous Treatment Effects by Combining Weak Instruments and Observational Data

### v1

@@ 5.2 Impact of 401(k) Participation on Financial Wealth
We demonstrate our method’s effectiveness with a real-world case study on the impact of 401(k) participation on financial wealth, using data from the 1991 Survey of Income and Program Participation [9]. The dataset includes 9,915 respondents with nine covariates: age, income, education, family size, marital status, two-earner status, pension status, IRA participation, and home ownership. The primary variable of interest is 401(k) participation ($A$), with eligibility ($Z$) as the instrumental variable. Although 401(k) eligibility is not randomly assigned, it is argued to maintain conditional independence given observed features [38, 9]. We assume 401(k) eligibility influences net worth only through 401(k) participation, characterizing this as an IV study with one-sided non-compliance, where non-eligible individuals cannot participate.. The target variable ($Y$) is net financial assets, calculated as the total of 401(k) balance, bank account balances, and interest-earning assets, minus non-mortgage debt.
@@
To replicate the scenario, we split the dataset into two halves: one for the IV study and the other for the observational study (where the instrument information is removed). Our goal is to use the parametric extension approach in Algorithm 1 to recover unbiased CATEs. Due to one-sided non-compliance, the estimated compliance factor $\widehat{\gamma}(x)$ is high ($0.49-0.90$, see Appendix C). We introduce non-compliance by setting $\gamma(x)$ to 0 for individuals with less than 12 years of education (13% of the population). In the first stage of Algorithm 1, we use RF regressors and classifiers to estimate $\tau^{O}(x)$, $\gamma(x)$, and $\pi_{Z}(x)$, with hyperparameters from [10]. In the second stage, we define the mapping $\phi(x)$ with an intercept term, the 9 covariates, and their interactions (46 features total). We apply mild $L_{1}$ regularization in the final linear regression due to the large number of features.
@@
In Figure 2, we analyze how the CATE function from Algorithm 1 varies with education. We focus on education because compliance selects education as a top importance feature, whereas the outcome models do not ((see Appendix C)) We vary education and marital status, setting age and income to the median and binary variables to zero. Since compliance in the IV study is high, we consider $\widehat{\tau}^{E}$ without artificial non-compliance as the ground truth. We observe that observational data treatment effects are upwardly biased due to unobserved confounders (e.g., financial literacy). The $\widehat{\tau}(x)$ from Algorithm 1, shown with a dashed line for non-compliance regions, closely follows $\widehat{\tau}^{E}(x)$. This demonstrates that combining IV and observational data effectively estimates unbiased CATEs in real-world scenarios, proving robust for causal inference with low compliance and unobserved confounding.
@@ 6 Conclusion
This study introduces a method that combines observational and instrumental variable (IV) data to tackle unobserved confounders in observational studies and low compliance in IV studies. Our two-stage framework first identifies biased CATEs from observational data, then corrects them using compliance-weighted IV samples. We explore two versions: one with a parametric form for confounding bias and another with a shared representation between true and biased CATE. We showed that both methods are consistent and validated through simulations and real-world applications. Our method is promising for applications in digital platforms, personalized medicine, and economics. This research represents an important advancement toward more reliable and actionable insights in complex data environments. Limitations of our work are discussed in Appendix D.
Figure 2: Impact of 401(k) participation on net worth by education level: Using $\widehat{\tau}$ from Algorithm 1, we fix age, income, and binary variables, varying education and marital status. The black line shows results from Algorithm 1, and the dashed line indicates predictions in the no-compliance region. $\widehat{\tau}^{O}$ is the observational CATE, while $\widehat{\tau}^{E}$ is the IV CATE without non-compliance.
@@ Appendix A Extended Literature Review
Heterogeneous treatment effect estimation from observational data
Recently, there has been a significant interest in applying machine learning to estimate CATEs using observational data. This field has seen adaptations of a wide range of machine learning techniques, from random forests [46, 34] and Bayesian algorithms e.g. [21, 17] to deep learning [24, 5, 43] and blackbox meta-learners [30, 33] that utilize efficient influence functions [28, 12] and Neyman orthogonality [10, 14]. Despite these advancements, a significant challenge persists as these methods typically assume the absence of confounding in observational data (ignorability) – an often unrealistic and unverifiable assumption – limiting their real-world applicability. Without ignorability, point identification of effects is impossible, although some studies propose methods to construct bounds on treatment effect estimates under assumptions about the structure of unobserved confounding [40, 25, 48, 19, 35]. Nonetheless, these bounds often have limited practical utility.
Heterogeneous treatment effect estimation using IVs
@@
Treatment effect estimation with weak instruments
@@
Combining observational and randomized data
There has been an proliferation of research in combining observational datasets with randomized control trials (RCT) – experimental data with perfect compliance – to mitigate bias from observational studies. One of the strategies is to impose structural assumptions such as strong parametric assumptions for the confounding bias [26] or a shared structure between the biased and unbiased CATE functions that can be estimated from the two datasets [20]. Other studies advocate for dual estimators from both data types, optimizing bias correction through a weighted average [50, 8, 41]. Additionally, approaches like those by [6] and [23] leverage outcomes from different time-steps, such as short-term and long-term effects, to enhance estimation accuracy. Our work is closest to [26] and [20]. However, our study faces additional complexities: firstly, the CATE estimation techniques differ between the datasets, requiring us to debias the overall effect function rather than just individual outcome functions. Secondly, while randomized control trials (RCTs) may not represent the target population due to their narrow scope, our instrumental variable (IV) study faces representation issues due to minimal or absent compliance in strata that are not known a priori. Thirdly, the CATE estimation in our IV study uses a ratio estimator, which is highly sensitive to changes in the compliance denominator, adding a layer of complexity to our analysis.
@@ Appendix B Proofs of Theorems and Lemmas
B.1 Proof of Lemma 1
Recall that for any $x\in\mathcal{X}^{\prime}$, we have that $\gamma(x)>0$ by Assumption 2. Then, assuming that $\pi_{Z}(x)>0$, we use the law of total expectation as follows:
@@ B.2 Proof of Theorem 2
We note that these quantities are well defined because $\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\right]$ is invertible. This is derived from the first and last conditions of Assumption 3, along with the stipulation in Assumption 1 that $\gamma(x)>0$ for all $x$ in a set of non-zero measure. Using these definitions, we can write
@@
$\leq\underbrace{\left\|\widehat{\theta}-\widetilde{\theta}_{n_{E}}\right\|_{2}}_{\lambda_{1}}+\underbrace{\left\|\widetilde{\theta}_{n_{E}}-\widetilde{\theta}\right\|_{2}}_{\lambda_{2}}+\underbrace{\left\|\widetilde{\theta}-\theta\right\|_{2}}_{\lambda_{3}}$ (Cauchy-Schwartz)
@@
$\qquad\cdot\mathbb{E}\left[\left(YZ(1-\pi_{Z}(X))-Y(1-Z)\pi_{Z}(X)+w(X)\tau(X)-w(X)\theta^{T}\phi(X)\right)w(X)\phi(X)\right]$ (Realizability of $b(X)$)
$=\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\mid\gamma(X)>0\right]^{-1}P(\gamma(X)>0)^{-1}$
$\qquad\cdot\Big{(}\mathbb{E}\left[\left(YZ(1-\pi_{Z}(X))-Y(1-Z)\pi_{Z}(X)-w(X)\tau(X)\right)w(X)\phi(X)\mid\gamma(X)>0\right]$
$\quad\qquad+\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\theta\mid\gamma(X)>0\right]\Big{)}P(\gamma(X)>0)$
$=\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\mid\gamma(X)>0\right]^{-1}$
$\qquad\cdot\mathbb{E}\left[\left(YZ(1-\pi_{Z}(X))-Y(1-Z)\pi_{Z}(X)-w(X)\tau(X)\right)w(X)\phi(X)\mid\gamma(X)>0\right]+\theta$
$=\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\mid\gamma(X)>0\right]^{-1}$
$\qquad\cdot\mathbb{E}\left[\left(\frac{YZ}{\pi_{Z}(X)\gamma(X)}-\frac{Y(1-Z)}{(1-\pi_{Z}(X))\gamma(X)}-\tau(X)\right)w^{2}(X)\phi(X)\Bigg{\lvert}\gamma(X)>0\right]+\theta$ (Since $\gamma(X)>0$ implies $w(X)>0$ by Assumption 3)
@@
By our boundedness assumptions, the second term yields an $O_{p}\left(\|w-\widehat{w}^{k}\|_{L_{2}}\right)=O_{p}(r_{\gamma}(n_{E})+r_{\pi_{Z}}(n_{E}))$ term in the expression for $O_{p}\left(\|\widehat{\Sigma}-\widehat{\Sigma}_{K}\|_{F}\right)$. To analyze the first term, let $E_{k}$ represent the samples in the $k^{\text{th}}$ fold of the $E$ dataset. Then, $\delta_{k}\mid E_{k}$ has mean 00 since $\widehat{w}^{(k)}$ is independent from $E_{k}$ due to the $K$-fold sample splitting. Then, we can apply Chebyshev’s inequality to obtain that
$\delta_{k}\mid E_{k}=O_{p}\left(n_{E}^{-1/2}\;\mathbb{E}\left[\left(w(X)^{2}-\widehat{w}^{(k)}(X)^{2}\right)^{2}\phi(X)^{2}_{i}\phi(X)^{2}_{j}\Big{\lvert}E_{k}\right]^{2}\right)=o_{p}(1/\sqrt{n_{E}})$
@@
Given that $\|\widehat{\tau}-\tau\|_{L_{2}}=\|(\theta-\widehat{\theta})^{T}\phi(X)+(\tau^{O}(X)-\tau^{O}(X))\|_{L_{2}}$, we further have
@@ B.3 Proof of Theorem 3
$=\Big{\|}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)\widehat{\phi}(X)^{T}\Big{\lvert}\gamma(X)>0\right]^{-1}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)\phi(X)^{T}\nu\Big{\lvert}\gamma(X)>0\right]-\nu\Big{\|}_{2}$ (Lemma 1)
$=\Big{\|}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)\widehat{\phi}(X)^{T}\Big{\lvert}\gamma(X)>0\right]^{-1}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)(\phi(X)-\widehat{\phi}(X))^{T}\nu\Big{\lvert}\gamma(X)>0\right]\Big{\|}_{2}$

### v2

@@ 5.2 Impact of 401(k) Participation on Financial Wealth
Figure 3: Impact of 401(k) participation on net worth by education level: Using $\widehat{\tau}(x)$ from Algorithm 1, we fix age, income, and binary variables, varying education and marital status. The black line shows results from Algorithm 1, and the dashed line indicates predictions in the no-compliance region. $\widehat{\tau}^{O}(x)$ is the biased observational CATE, while $\widehat{\tau}^{E}(x)$ is the IV CATE without non-compliance.
@@
We demonstrate our method’s effectiveness with a real-world case study on the impact of 401(k) participation on financial wealth, using data from the 1991 Survey of Income and Program Participation [11]. The dataset includes 9,915 respondents with nine covariates: age, income, education, family size, marital status, two-earner status, pension status, IRA participation, and home ownership. The primary variable of interest is 401(k) participation ($A$), with eligibility ($Z$) as the instrumental variable. Although 401(k) eligibility is not randomly assigned, it is argued to maintain conditional independence given observed features [43, 11]. We assume 401(k) eligibility influences net worth only through 401(k) participation, characterizing this as an IV study with one-sided non-compliance, where non-eligible individuals cannot participate ($A^{E}(0)=0$). The target variable ($Y$) is net financial assets, calculated as the total of 401(k) balance, bank account balances, and interest-earning assets, minus non-mortgage debt.
@@
To replicate the scenario in this paper, we split the dataset into two halves: one for the IV study and the other for the observational study (where we intentionally remove the instrument information). Our goal is to use these datasets, along with the parametric extension approach in Algorithm 1, to recover the unbiased CATEs. Due to one-sided non-compliance, the estimated compliance factor $\widehat{\gamma}(x)$ is high ($0.49-0.90$, see Appendix C). To show the utility of our method, we introduce artificial non-compliance by setting $\gamma(x)$ to 0 for individuals with less than 12 years of education (13% of the population). In the first stage of Algorithm 1, we use RF regressors and classifiers to estimate $\tau^{O}(x)$, $\gamma(x)$, and $\pi_{Z}(x)$, with hyperparameters set based on other related work on this dataset [12]. In the second stage, we define the mapping $\phi(x)$ with an intercept term, the 9 covariates, and their interactions (46 features total). We apply a mild $L_{1}$ regularization in the final linear regression due to the large number of resulting features.
@@
In Figure 3, we study how the CATE function from Algorithm 1 varies with education. We focus on education as it is selected as a top feature by the compliance model, while the outcome models do not rank it as highly significant (see Appendix C). To explore this relationship, we vary education and marital status, holding age and income at their median values and setting all binary variables to zero. Since compliance in the IV study is high, we consider the estimate $\widehat{\tau}^{E}(x)$ without the artificial non-compliance as the ground truth for comparison. Our analysis shows that observational data treatment effects are upwardly biased, likely due to unobserved confounders such as financial literacy. The $\widehat{\tau}(x)$ from Algorithm 1, shown with a dashed line for non-compliance regions, closely aligns with $\widehat{\tau}^{E}(x)$ (excluding the artificial non-compliance). This demonstrates that combining IV and observational data can effectively estimate unbiased CATEs in real-world settings, offering a robust solution for causal inference even in the presence of low compliance and unobserved confounding.
@@ 6 Conclusion
This study introduces a method that combines observational and instrumental variable (IV) data to address unobserved confounders in observational studies and low compliance in IV studies. Our two-stage framework first estimates biased CATEs from the observational data, then corrects them using compliance-weighted IV samples. We explore two variations of our procedure: one that models confounding bias parametrically, and another that leverages a shared representation between the true and biased CATEs. Both methods are shown to be consistent, validated through simulations and real-world applications. Our approach holds significant promise for applications in digital platforms, personalized medicine, and economics, offering a robust tool for deriving reliable, actionable insights from complex data. Limitations of our work are discussed in Appendix D.
Acknowledgements
We thank the anonymous reviewers for their valuable feedback and insightful suggestions. This material is based upon work supported by the National Science Foundation under Grant No. 1846210 and by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0023112.
@@ Appendix A Extended Literature Review
Heterogeneous treatment effect estimation from observational data:
Recently, there has been a significant interest in applying machine learning to estimate CATEs using observational data. This field has seen adaptations of a wide range of machine learning techniques, from random forests [51, 39] and Bayesian algorithms e.g. [24, 20] to deep learning [27, 5, 48] and blackbox meta-learners [33, 38] that utilize efficient influence functions [31, 14] and Neyman orthogonality [12, 16]. Despite these advancements, a significant challenge remains as these methods typically assume the absence of confounding in observational data (ignorability) – an often unrealistic and unverifiable assumption – limiting their real-world applicability. Without ignorability, point identification of effects is impossible, although some studies propose methods to construct bounds on treatment effect estimates under assumptions about the structure of unobserved confounding [45, 28, 40, 17]. Nonetheless, these bounds often have limited practical utility. Other efforts to address confounding bias in CATE estimation rely on latent variable models to recover unobserved confounders from noisy proxies [36, 34] or utilize multiple or sequential treatments [53, 9, 22]. However, these methods also have limited practical impact, as they depend on either the availability of additional accurate proxy data or unverifiable assumptions such as no unobserved single-cause confounders.
Heterogeneous treatment effect estimation using IVs:
@@
Treatment effect estimation with weak instruments:
@@
Combining observational and randomized data:
There has been a proliferation of research in combining observational datasets with randomized control trials – experimental data with perfect compliance – to mitigate bias from observational studies. One of the strategies is to impose structural assumptions such as strong parametric assumptions for the confounding bias [29] or a shared structure between the biased and unbiased CATE functions that can be estimated from the two datasets [23]. Other studies advocate for dual estimators from both data types, optimizing bias correction through a weighted average [55, 10, 46]. Additionally, approaches like those by [6] and [26] leverage outcomes from different time-steps, such as short-term and long-term effects, to enhance estimation accuracy. Our work is closest to [29] and [23]. However, our study faces additional complexities: firstly, the CATE estimation techniques differ between the datasets, requiring us to debias the overall effect function rather than just individual outcome functions. Secondly, RCTs may not represent the target population due to their narrow scope, our instrumental variable (IV) study faces representation issues due to minimal or absent compliance in strata that are not known a priori. Thirdly, the CATE estimation in our IV study uses a ratio estimator, which is highly sensitive to changes in the compliance denominator, adding a layer of complexity to our analysis.
@@ Appendix B Proofs of Theorems and Lemmas
B.1 Proof of Equation 3
The exclusion and independence conditions in Assumption 1 imply that the following identification equation holds:
$\mathbb{E}\left[Y^{E}\left(A^{E}(1)\right)-Y^{E}\left(A^{E}(0)\right)\mid X^{E}=x\right]$ (8)
$\qquad=\mathbb{E}[Y^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[Y^{E}\mid Z^{E}=0,X^{E}=x].$
By noting that
$\begin{cases}Y^{E}\left(A^{E}(1)\right)-Y^{E}\left(A^{E}(0)\right)=Y(1)-Y(0),&\text{when }A^{E}(1)=1,A^{E}(0)=0\;(\text{compliers})\\ Y^{E}\left(A^{E}(1)\right)-Y^{E}\left(A^{E}(0)\right)=Y(0)-Y(1),&\text{when }A^{E}(1)=0,A^{E}(0)=1\;(\text{defiers})\\ Y^{E}\left(A^{E}(1)\right)-Y^{E}\left(A^{E}(0)\right)=0,&\text{when }A^{E}(1)=1,A^{E}(0)=1\;(\text{always-takers})\\ Y^{E}\left(A^{E}(1)\right)-Y^{E}\left(A^{E}(0)\right)=0,&\text{when }A^{E}(1)=0,A^{E}(0)=0\;(\text{never-takers})\\ \end{cases}$
the left-hand side of this equation can further be written as:
$\mathbb{E}[Y^{E}(A^{E}(1))-Y^{E}(A^{E}(0))\mid X^{E}=x]$ (9)
$=\mathbb{E}[(Y^{E}(1)-Y^{E}(0))(A^{E}(1)-A^{E}(0))\mid X^{E}=x]$
$=\mathbb{E}[Y^{E}(1)-Y^{E}(0)\mid X^{E}=x]\cdot\mathbb{E}[A^{E}(1)-A^{E}(0)\mid X^{E}=x]$ (Assumption 2)
$=\tau(x)\cdot(\mathbb{E}[A^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[A^{E}\mid Z^{E}=0,X^{E}=x])$ (Assumption 1)
Since the claim of Equation 3 holds for $x\in\mathcal{X}^{\prime}$, we have that $\mathbb{E}[A^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[A^{E}\mid Z^{E}=0,X^{E}=x]\neq 0$ by the relevance condition in Assumption 1. From Eqs. 8 and 9, we obtain:
$\tau(x)$ $=\frac{\mathbb{E}[Y^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[Y^{E}\mid Z^{E}=0,X^{E}=x]}{\mathbb{E}[A^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[A^{E}\mid Z^{E}=0,X^{E}=x]}$
for $x\in\mathcal{X}^{\prime}$.
B.2 Proof of Lemma 1
Recall that for any $x\in\mathcal{X}^{\prime}$, we have that $\gamma(x)\neq 0$ by Assumption 2. Then, assuming that $\pi_{Z}(x)>0$, we use the law of total expectation as follows:
@@ B.3 Proof of Theorem 2
We note that these quantities are well defined because $\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\right]$ is invertible. This follows from the first and last conditions of Assumption 3, along with the stipulation in Assumption 1 that $\gamma(x)\neq 0$ for all $x$ in a set of non-zero measure. Using these definitions, we can write
@@
$\leq\underbrace{\left\|\widehat{\theta}-\widetilde{\theta}_{n_{E}}\right\|_{2}}_{\lambda_{1}}+\underbrace{\left\|\widetilde{\theta}_{n_{E}}-\widetilde{\theta}\right\|_{2}}_{\lambda_{2}}+\underbrace{\left\|\widetilde{\theta}-\theta\right\|_{2}}_{\lambda_{3}}$ (Triangle Inequality)
@@
$\qquad\cdot\mathbb{E}\left[\left(YZ(1-\pi_{Z}(X))-Y(1-Z)\pi_{Z}(X)-w(X)\tau(X)+w(X)\theta^{T}\phi(X)\right)w(X)\phi(X)\right]$ (Realizability of $b(X)$)
$=\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\mid\gamma(X)\neq 0\right]^{-1}P(\gamma(X)\neq 0)^{-1}$
$\qquad\cdot\Big{(}\mathbb{E}\left[\left(YZ(1-\pi_{Z}(X))-Y(1-Z)\pi_{Z}(X)-w(X)\tau(X)\right)w(X)\phi(X)\mid\gamma(X)\neq 0\right]$
$\quad\qquad+\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\theta\mid\gamma(X)\neq 0\right]\Big{)}P(\gamma(X)\neq 0)$
$=\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\mid\gamma(X)\neq 0\right]^{-1}$
$\qquad\cdot\mathbb{E}\left[\left(YZ(1-\pi_{Z}(X))-Y(1-Z)\pi_{Z}(X)-w(X)\tau(X)\right)w(X)\phi(X)\mid\gamma(X)\neq 0\right]+\theta$
$=\mathbb{E}\left[w(X)^{2}\phi(X)\phi(X)^{T}\mid\gamma(X)\neq 0\right]^{-1}$
$\qquad\cdot\mathbb{E}\left[\left(\frac{YZ}{\pi_{Z}(X)\gamma(X)}-\frac{Y(1-Z)}{(1-\pi_{Z}(X))\gamma(X)}-\tau(X)\right)w(X)^{2}\phi(X)\Bigg{\lvert}\gamma(X)\neq 0\right]+\theta$ (Since $\gamma(X)\neq 0$ implies $w(X)\neq 0$ by Assumption 3)
@@
By our boundedness assumptions, the second term yields an $O_{p}\left(\|w-\widehat{w}^{k}\|_{L_{2}}\right)=O_{p}(r_{\gamma}(n_{E})+r_{\pi_{Z}}(n_{E}))$ term in the expression for $O_{p}\left(\|\widehat{\Sigma}-\widehat{\Sigma}_{K}\|_{F}\right)$. To analyze the first term, let $E_{k}$ represent the samples in the $k^{\text{th}}$ fold of the $E$ dataset. Then, $\delta_{k}\mid E_{k}$ has mean 00 since $\widehat{w}^{(k)}$ is independent from $E_{k}$ due to the $K$-fold sample splitting. Then, we can apply Chebyshev’s inequality to obtain
$\delta_{k}\mid E_{k}=O_{p}\left(n_{E}^{-1/2}\;\mathbb{E}\left[\left(w(X)^{2}-\widehat{w}^{(k)}(X)^{2}\right)^{2}\phi(X)^{2}_{i}\phi(X)^{2}_{j}\Big{\lvert}E_{k}\right]^{1/2}\right)=o_{p}(1/\sqrt{n_{E}})$
@@
Given that $\|\widehat{\tau}-\tau\|_{L_{2}}=\|(\theta-\widehat{\theta})^{T}\phi(X)+(\tau^{O}(X)-\widehat{\tau}^{O}(X))\|_{L_{2}}$, we further have
@@ B.4 Proof of Theorem 3
$=\Big{\|}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)\widehat{\phi}(X)^{T}\Big{\lvert}\gamma(X)\neq 0\right]^{-1}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)\phi(X)^{T}\nu\Big{\lvert}\gamma(X)\neq 0\right]-\nu\Big{\|}_{2}$ (Lemma 1)
$=\Big{\|}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)\widehat{\phi}(X)^{T}\Big{\lvert}\gamma(X)\neq 0\right]^{-1}\mathbb{E}\left[w(X)^{2}\widehat{\phi}(X)(\phi(X)-\widehat{\phi}(X))^{T}\nu\Big{\lvert}\gamma(X)\neq 0\right]\Big{\|}_{2}$